Main INDEX
Monthly INDEX
PREV
NEXT
User name Adaq
Log entry time 08:11:43 on June 8,2005
Entry number 145507
keyword=Shift Summary
Running at 50 uA, will continue to do this unitl 2am
finished run 4394 at 00:23
00:24 start run 4395
00:51 start run 4394
01:23 start run 4397
01:50 start run 4398
02:20 switch to 100 uA
02:22 start run 4399
02:34 start run 4400
02:50 start and stopped run 4401, lost beam, many trips
go back to 52 uA
3:55 start run 4402
4:17 start run 4403
4:45 start run 4404
beam becomes 'useless' -> stop run 4405
get stable 60uA beam
6:51 start 4406
7:10 MCC needed the beam to study the optics
Beam trips : about 20 in 2hrs @ 50uA
about 10 in 20min @100 uA (at 2:50), no stable beam
possible, many ionchamber trips
increasing number of beam trips after 05:00, beam
unstable even at 50uA